If 1 of alloy A is 10%, and 1 of alloy B is 50%, how many of each would you need to combine together to get 10 at 40%?
Answer by jorel1380(3719)

x+y=10
.1x+.5y=4
x+5y=40
4y=30
y=30/4
x=10/4
You need 10/4ths of 10% alloy, and 30/4ths of 50% alloy..